OpenStudy (anonymous):

Can the imaginary number i be raised to the power of a fractional number, say 1/2 or 1/4?

OpenStudy (solomonzelman):

in theory, why not?

OpenStudy (anonymous):

Please elaborate. How would I be able to do this?

OpenStudy (anonymous):

understand that \[b^x=e^{x\ln(b)}\] would help

OpenStudy (anonymous):

the answer is not unique for complex numbers \[i^{\frac{1}{2}}=\sqrt{i}\] or a number whose square is \(i\) there are two such numbers